CH/Doc/92/03

Charissa

List of references

Author: Wilton Catford Date: August 13, 1992

Cumulative List of References

Source copies of these documents are kept on file at Surrey. If you require copies, please contact W.N. Catford (wnc@dl.nnga).

Table 1: Minutes of Charissa Group Meetings

Ref. No.  Venue  Meeting Date Minutes Date
CH/Meet/92/01  Birmingham 4 March 92  5 March 92
CH/Meet/92/02  Daresbury  29 April 92  8 May 92
CH/Meet/92/03  Oxford  11 June 92  13 Aug 92

Table 2: Agendas for Charissa Group Meetings

Ref. No.  Venue  Meeting Date
CH/Agen/92/01  Birmingham 4 March 92
CH/Agen/92/02  Daresbury  29 April 92
CH/Agen/92/03  Oxford  11 June 92
CH/Agen/92/04  Birmingham 24 August 92

Table 3: Technical Reports

Ref. No.  Author  Title  Date
CH/Tech/92/01  B.R. Fulton  Silicon Detector News  27 April 92
CH/Tech/92/02  N.M. Clarke  Progress on Scintillation Detectors  27 Feb 92
CH/Tech/92/03  N.M. Clarke  Letter re Fitting Code  16 March 92
CH/Tech/92/04  R.A. Hunt  Proposals for Charissa Electronics  28 April 92
CH/Tech/92/05  N.M. Clarke  Progress on Scintillation Detectors  8 May 92
CH/Tech/92/06  B.R. Fulton  Computing, Data Acquisition and Control  19 May 92
CH/Tech/92/07  B.R. Fulton  Silicon Detector Developments  10 June 92
CH/Tech/92/08  N.M. Clarke  Test on Prototype CsI Scintillator  11 June 92

CH/Tech/92/09 W.D.M. Rae/D.Smith New Bias Supply for Charissa 19 May 92

CH/Tech/92/10  R.A. Hunt  Charissa Amplifier Development  10 June 92

Table 4: Documents and Non-technical Reports

Ref. No.  Author  Title  Date

CH/DOC/92/01 W.N. Catford List of Names of Collaborators 13 Aug 92 CH/DOC/92/02 N.M. Clarke Update on the ICARE project at

Strasbourg

28 March 92

CH/DOC/92/03 W.N. Catford This cumulative list August 13, 1992 CH/DOC/92/04 W.D.M. Rae Matchstick Calibrations 2 June 92 CH/DOC/92/05 B.F. Fulton Proposal for a Joint French–UK

collaboration 10 April 92